OpenStudy (anonymous):

14=4a-5 i got a=5 is this correct?

OpenStudy (jackellyn):

if you want to check your answer substitute 5 where you see an a in the equation 14=4(5)-5 If both sides are 14 then that is your answer. So, is a=5 your answer?

jimthompson5910 (jim_thompson5910):

Side note: 4(5) means 4 times 5

OpenStudy (waleed_imtiaz):

It wull give 14=15.... thats not true..... a=19/4.. then 14=14

OpenStudy (jackellyn):

thanks @jim_thompson5910 for the backup :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jackellyn i don't understand what your saying

jimthompson5910 (jim_thompson5910):

Take 14=4a-5 and replace 'a' with 5 (since you claim that a = 5) to get 14=4(5)-5 This is the same as saying "14 = 4 times 5 - 5"

OpenStudy (anonymous):

oh ok thank you

jimthompson5910 (jim_thompson5910):

If you can show that the right side is equal to the left, then a = 5 is a solution. If not, then it's not a solution.

OpenStudy (anonymous):

ok